[figure: 9.7 cm]

Magnetostriction and magnetization oscillations of superconducting

2H-NbSe2 single crystals

V.V. Eremenko^{a}, \underline{Yu.A. Shabakayeva}^{a}, V.A. Sirenko^{a}, R. Schleser^{b} aInstitute for Low Temperature Physics and Engineering, 61103 Kharkov, Ukraine

bGrenoble High Magnetic Field Laboratory, MPI-FKF and CNRS BP-166, F-38042, Grenoble Cedex 09, France

We present magnetostriction and magnetization measurements on 2H- NbSe2 single crystals in magnetic field up to 20 T and at temperature between 1.5 K and 8K. Using the amplitudes of the oscillatory magnetostriction component and of the oscillatory magnetization component we get the important parameter ∂lnS_{m}/∂σ_{α}, where S_{m} is the extremal cross-sectional area of the Fermi surface, σ_{α} is the mechanical stress, and, therefore, that gave us detailed information on the response of the Fermi surface to mechanical stress. The expression for ∂lnS_{m}/∂σ_{α} was obtained from a thermodynamic analysis of the oscillating part of the thermodynamic potential. The values obtained are much higher than the values of ∂lnS_{m}/∂σ_{α} for common metals. This is due to the peculiar geometry of Fermi surface in this anisotropic compound. The periods of magnetostriction and magnetization oscillations coincide, which means that both effects are attributed to Landau quantum oscillations. The periods, frequencies and amplitudes of the oscillations can be used to extract all the information about electronic structure (e.g. effective masses, shape and size of the Fermi surface) of this compound.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : Yu.Shabakayeva

Address for correspondence : Yu.Shabakayeva Institute for Low Temperature Physics and Engineering 61103 Kharkov, Ukraine

Email address : shabakayeva@ilt.kharkov.ua